# LLaSE-G1: Incentivizing Generalization Capability for LLaMA-based Speech Enhancement

### v2

@@ 3.1 Overall Architecture
Formally, let: 1. $\texttt{Vertorize}(X)=\{x_{1},\dots,x_{N}\}$ be the WavLM encoder, which converts input degraded speech $X$ into $N$ speech features. 2. $\texttt{Vertorize}(P)=\{p_{1},\dots,p_{T}\}$ be the WavLM encoder, which converts optional reference speech $X$ into $T$ speech features. 3. $\texttt{Tokenize}(Y)=\{y_{1},\dots,y_{M}\}$ be the X-codec2 encoder, which converts an enhanced speech $Y$ into $M$ speech tokens. 4. $\texttt{Detokenize}(\{y_{1},\dots,y_{M}\})=\hat{Y}$ be the X-codec2 decoder, which reconstructs the waveform $\hat{Y}$ from its token representations. As the downsampling rate of WavLM is the same as that of X-codec2, $N$ is equal to $M$. Given a dataset $\mathcal{D}=\{(X_{i},P_{i},Y_{i})\}_{i=1}^{S}$, where $X_{i}$ is the degraded speech, $Y_{i}$ is the enhanced speech and $P_{i}$ is the reference speech or empty if unavailable, we represent each pair $(X_{i},P_{i},Y_{i})$ as a sequence $(x_{1},\dots,x_{N},p_{1},\dots,p_{T},y_{1},\dots,y_{M})$. Since the $X_{i}$ and $P_{i}$ are always given as input during training and inference, we pad $X_{i}$ and $P_{i}$ to the same length and the LM $\theta$ focuses on learning the conditional probability:
@@ 3.3 Unifying Various SE tasks
While NS, PLC, AEC, and TSE tasks output a single enhanced speech, we note that AEC requires the removal of information related to the reference speech, while TSE necessitates the preservation of reference speech information. To address this, we introduce a dual-channel output with two linear projection heads to unify the output representations. The first channel $c_{1}$ predicts tokens related to reference speech and the second channel $c_{2}$ predicts tokens irrelevant to reference speech. With these designs, for tasks including NS, AEC, and PLC, we employ a single-supervision strategy $\mathcal{L}_{\text{S}}$ through the cross-entropy loss between $c_{0}$ and the tokens $t_{0}$ extracted from the clean signal:
@@ 4.1 Experimental Setup
Data augmentation. We utilized dynamic data augmentation during training. For the NS task, the clean audio and noise are randomly selected and mixed with a signal-to-noise Ratio (SNR) ranging from [-5,20] dB. Both clean and noisy signals have a 50% probability of adding reverberation. In the PLC task, we use a two-state first-order Markov chain to describe the packet loss status of the current frame and the next frame. The transition and hold probabilities for Markov states are selected between 0.05 and 0.95. We directly generate a binary mask sequence and apply it to the clean speech. For the AEC task, we randomly select a real echo signal and its corresponding reference signal from the far-end single talk in the AEC Challenge dataset. The signal-to-echo ratio (SER) ranges from -15 dB to 15 dB. Noise is added with a 20% chance, and the SNR is between -5 dB and 20 dB. For the Target Speech Enhancement (TSE) task, we select a clean speech segment and its corresponding auxiliary segment for the enrollment speech, while a different speaker is chosen for the interference speech. There is a 5% probability that no interfering speaker is present. Target speech and interfering speech are mixed with an SNR ranging from [-15, 15] dB, with an additional 10% probability of adding extra noise.
@@
Evaluation Metrics. We use objective metrics to evaluate the performance of the baseline systems and our model. DNSMOS Reddy et al. (2022) include speech quality (SIG), background noise quality (BAK), and overall quality (OVRL) of the audio. AECMOS Purin et al. (2022) consists of echo annoyance MOS (EMOS) and other degradation MOS (DMOS). PLCMOS Diener et al. (2023) is used to assess the quality of audio processed by PLC algorithms. All MOS scores range from 1 to 5, representing audio quality from low to high. SpeechBERTScore (SBS) Saeki et al. (2024) is also employed to evaluate the semantic similarity between the enhanced audio and the reference audio. Following Zhang et al. (2025), we use HuBERT-base [^4] Hubert-base on Hugging Face model to extract semantic features. For acoustic similarity, we calculate speaker similarity SimWB based on the WavLM-base-sv model [^5] WavLM-base-sv on Huggingface to evaluate the performance.
@@ 4.2.1 Noise Suppression
Table 2 presents a comparison between the proposed LLaSE-G1 and several SOTA discriminative and generative models. The "With Reverb" column represents the test set with reverberation, and the "No Reverb" column is the one without. The results indicate that generative NS models consistently outperform discriminative models, especially in reverberant conditions. With single inference, LLaSE-G1 already surpassed most other systems. After multiple inferences, its performance improves further, achieving a SOTA result of 3.49 OVRL score at the no_reverb test set and 3.42 OVRL score at the with_reverb test set.
@@ 4.2.3 Target Speaker Extraction
As shown in Table 4, in both tracks, LLaSE-G1’s SIG MOS significantly outperformed other methods, indicating that LMs-based generative models provide higher audio quality with less signal distortion. The OVRL scores suggest that TEA-PSE 3.0 and NAPSE have advantages on headset and speakerphone devices, respectively. However, the proposed LLaSE-G1 achieved the best performance across both tracks, demonstrating superior device generalization capabilities compared to discriminative models.
@@ 4.2.6 Semantic and Speaker Similarity
Table 7: Semantic and speaker similarity results on various tasks, using the same test sets from previous subsections.
@@ 4.2.7 Ablation Study
We conduct an ablation study to evaluate the effectiveness of input representations, output representations, and model backbone, choosing SELM as the baseline. As shown in Table 8, when replacing inputs and output with proposed continuous features and speech tokens, there is an obvious improvement, revealing the effectiveness of acoustic preservation. Besides, there is no performance drop when replacing the full attention Transformer with casual attention LLaMA. Finally, adopting a multi-inference strategy further boosts performance
Table 8: DNSMOS scores on DNS blind test set without reverb. "D" represents Discrete tokens, and "C" represents Continuous features. "S" represents Single inference, and "M" represents Multiple inference.
@@
Baseline D Transformer HiFiGAN S 3.26
C Transformer HiFiGAN S 3.34
C LLaMA HiFiGAN S 3.35
Proposed C LLaMA X-codec2 S 3.43
C LLaMA X-codec2 M 3.49
@@ 5 Conclusion
In this study, we propose LLaSE-G1, a general LLaMA-based framework to unify various speech enhancement tasks. Specifically, we employ continuous features as input and predict 1D speech tokens, maximizing acoustic preservation. Additionally, we design dual-channel inputs and outputs, unifying multiple SE tasks. Extensive experiments show that LLaSE-G1 achieves superior performance in each benchmark, serving as a powerful foundation model. Moreover, LLaSE-G1 demonstrates scaling effects at test time and emerging capabilities for unseen SE tasks.
@@ A.1 Model Configuration
We use the open-source checkpoints of WavLM-large and X-codec2. The LLaMA-based LM comprises 16 LLaMA layers, each with 16 attention heads, a dropout rate of 0.1, a hidden size of 2048, and an intermediate size of 4096. The total number of parameters in the model is approximately 1.07 billion, which includes all learnable weights and biases across all layers. The model has 2 input linear layers and 2 output linear layers. The input layer maps a 1024-dimensional vector to another 1024-dimensional vector, while the output layer transforms a 2048-dimensional vector into a 65536-dimensional vector, which is the codebook size of Xcodec2.

### v4

@@ 3.1 Overall Architecture
Formally, let: 1. $\texttt{Vertorize}(X)=\{x_{1},\dots,x_{N}\}$ be the WavLM encoder, which converts input degraded speech $X$ into $N$ speech features. 2. $\texttt{Vertorize}(P)=\{p_{1},\dots,p_{T}\}$ be the WavLM encoder, which converts optional reference speech $P$ into $T$ speech features. 3. $\texttt{Tokenize}(Y)=\{y_{1},\dots,y_{M}\}$ be the X-codec2 encoder, which converts an enhanced speech $Y$ into $M$ speech tokens. 4. $\texttt{Detokenize}(\{y_{1},\dots,y_{M}\})=\hat{Y}$ be the X-codec2 decoder, which reconstructs the waveform $\hat{Y}$ from its token representations. As the downsampling rate of WavLM is the same as that of X-codec2, $N$ is equal to $M$. Given a dataset $\mathcal{D}=\{(X_{i},P_{i},Y_{i})\}_{i=1}^{S}$, where $X_{i}$ is the degraded speech, $Y_{i}$ is the enhanced speech and $P_{i}$ is the reference speech or empty if unavailable, we represent each pair $(X_{i},P_{i},Y_{i})$ as a sequence $(x_{1},\dots,x_{N},p_{1},\dots,p_{T},y_{1},\dots,y_{M})$. Since the $X_{i}$ and $P_{i}$ are always given as input during training and inference, we pad $X_{i}$ and $P_{i}$ to the same length and the LM $\theta$ focuses on learning the conditional probability:
@@ 3.3 Unifying Various SE tasks
While NS, PLC, AEC, and TSE tasks output a single enhanced speech, we note that AEC requires the removal of information related to the reference speech, while TSE necessitates the preservation of reference speech information. To address this, we introduce a dual-channel output with two linear projection heads to unify the output representations. The output embedding of the LLM is passed through these two linear projection to produce two embeddings from it. The first channel $c_{0}$ predicts tokens related to reference speech and the second channel $c_{1}$ predicts tokens irrelevant to reference speech. With these designs, for tasks including NS, AEC, and PLC, we employ a single-supervision strategy $\mathcal{L}_{\text{S}}$ through the cross-entropy loss between $c_{0}$ and the tokens $t_{0}$ extracted from the clean signal:
@@ 4.1 Experimental Setup
Data augmentation. We utilized dynamic data augmentation during training. For the NS task, the clean audio and noise are randomly selected and mixed with a signal-to-noise Ratio (SNR) ranging from [-5,20] dB. Both clean and noisy signals have a 50% probability of adding reverberation. In the PLC task, we use a two-state first-order Markov chain to describe the packet loss status of the current frame and the next frame. The transition and hold probabilities for Markov states are selected between 0.05 and 0.95. We directly generate a binary mask sequence and apply it to the clean speech. For the AEC task, we randomly select a real echo signal and its corresponding reference signal from the far-end single talk in the AEC Challenge dataset. The signal-to-echo ratio (SER) ranges from -15 dB to 15 dB. Noise is added with a 20% chance, and the SNR is between -5 dB and 20 dB. For the Target Speaker Extraction (TSE) task, we select a clean speech segment and its corresponding auxiliary segment for the enrollment speech, while a different speaker is chosen for the interference speech. There is a 5% probability that no interfering speaker is present. Target speech and interfering speech are mixed with an SNR ranging from [-15, 15] dB, with an additional 10% probability of adding extra noise.
@@
Evaluation Metrics. We use objective metrics to evaluate the performance of the baseline systems and our model. DNSMOS Reddy et al. (2022) include speech quality (SIG), background noise quality (BAK), and overall quality (OVRL) of the audio. AECMOS Purin et al. (2022) consists of echo annoyance MOS (EMOS) and other degradation MOS (DMOS). PLCMOS Diener et al. (2023) is used to assess the quality of audio processed by PLC algorithms. All MOS scores range from 1 to 5, representing audio quality from low to high. SpeechBERTScore (SBS) Saeki et al. (2024) is also employed to evaluate the semantic similarity between the enhanced audio and the reference audio. Following Zhang et al. (2025), we use HuBERT-base [^4] Hubert-base on Hugging Face model to extract semantic features. For acoustic similarity, we calculate speaker similarity SimWB based on the WavLM-base-sv model [^5] WavLM-base-sv on Huggingface to evaluate the performance. Subjective evaluations are also conducted using the Mean Opinion Score (MOS) as the primary metric to assess the model’s performance.
@@ 4.2.1 Noise Suppression
Table 2 presents a comparison between the proposed LLaSE-G1 and several state-of-the-art (SOTA) discriminative and generative models. The "With Reverb" column corresponds to the test set containing reverberation, while the "No Reverb" column refers to the clean test set without reverberation. The results show that generative noise suppression (NS) models consistently outperform discriminative ones, particularly under reverberant conditions. Even with single inference, LLaSE-G1 surpasses most competing systems. When employing multiple inferences, its performance improves further, achieving a SOTA OVRL score of 3.49 on the "No Reverb" test set and 3.42 on the "With reverb" test set.
Notably, LLaSE-G1 demonstrates exceptional performance on the with reverb test set compared to other generative enhancement systems, further highlighting the efficiency of continuous representations and the Xcodec2 decoder in handling challenging noise suppression tasks.
@@ 4.2.3 Target Speaker Extraction
As shown in Table 4, LLaSE-G1 consistently achieves significantly higher SIG MOS scores across both tracks, surpassing all other methods. This indicates that language model-based generative approaches offer higher audio quality with reduced signal distortion. While TEA-PSE 3.0 and NAPSE exhibit certain advantages on headset and speakerphone devices respectively, LLaSE-G1 delivers the best overall performance across both tracks, demonstrating superior device generalization compared to discriminative models.
@@ 4.2.6 Semantic and Speaker Similarity
Table 7: Semantic and speaker similarity results on various tasks, using the same test sets from previous subsections. As for the NS task, we report the average result across the No Reverb and With Reverb test sets.
@@
4.2.7 Subjective Evaluation
We also conducted evaluations on subjective listening tests and user studies. The experimental setup and results are shown in Table 8. We primarily used the DNS blind test set for the subjective evaluation. A total of 12 participants were recruited, and each participant was asked to listen to 50 audio samples. The Mean Opinion Score (MOS) was calculated on a 5-point scale, ranging from 1 (poor quality) to 5 (excellent quality). Although GENSE achieves comparable performance to LLaSE-G1 on the DNSMOS metric, LLaSE-G1 still delivers better perceptual quality in subjective listening, further demonstrating its superior capability in preserving acoustic details.
Table 8: Subjective MOS comparison. “D” denotes discriminative models, “G” denotes generative models.
Model Type MOS
Clean - 4.8125
FRCRN D 3.8917
SELM G 3.7750
GENSE G 4.0083
LLaSE-G1 G 4.1208
@@ 4.2.8 Ablation Study
We conduct an ablation study to evaluate the effectiveness of input representations, output representations, and model backbone, choosing SELM as the baseline. As shown in Table 9, when replacing inputs and output with proposed continuous features and speech tokens, there is an obvious improvement, revealing the effectiveness of acoustic preservation. Additionally, Experimental results show that while Xcodec2 tokens are designed to better capture acoustic information, discrete tokens still discard more detailed content than continuous embeddings when used as model inputs. Besides, there is no performance drop when replacing the full attention Transformer with casual attention LLaMA layers. Finally, adopting a multi-inference strategy further boosts performance.
Table 9: DNSMOS scores on DNS blind test set without reverb. "T" represents Discrete tokens, and "E" represents Embeddings. "S" represents Single inference, and "M" represents Multiple inference.
@@
Baseline WavLM Token Transformer HiFiGAN S 3.26
WavLM Embedding Transformer HiFiGAN S 3.34
WavLM Embedding LLaMA HiFiGAN S 3.35
Xcodec2 Token LLaMA X-codec2 S 3.29
Proposed WavLM Embedding LLaMA X-codec2 S 3.43
WavLM Embedding LLaMA X-codec2 M 3.49
@@ 5 Conclusion
In this study, we propose LLaSE-G1, a general LLaMA-based framework that unifies a wide range of speech enhancement (SE) tasks. Specifically, we utilize continuous acoustic features as input and predict 1D speech tokens to maximize fidelity to the original audio. To support multiple SE tasks, we design dual-channel inputs and outputs within a unified architecture. Extensive experiments demonstrate that LLaSE-G1 achieves state-of-the-art performance across various benchmarks, establishing it as a strong foundation model. Furthermore, LLaSE-G1 exhibits clear test-time scaling behavior and emerging generalization capabilities to previously unseen SE tasks. Notably, our framework requires no additional prompts to differentiate between SE tasks, instead relying on the LLM’s intrinsic ability to infer the task type from the input itself.
@@ A.1 Model Configuration
We use the open-source checkpoints of WavLM-large and X-codec2. The LLaMA-based LM comprises 16 LLaMA layers, each with 16 attention heads, a dropout rate of 0.1, a hidden size of 2048, and an intermediate size of 4096. The total number of parameters in the model is approximately 1.07 billion, which includes all learnable weights and biases across all layers. The model has 2 input linear layers and 2 output linear layers. The input layer maps the 1024-dimensional WavLM embedding to another 1024-dimensional vector, while the output layer transforms a 2048-dimensional vector into a 65536-dimensional vector, which is the codebook size of Xcodec2.